Suppose that systolic blood pressure for 35- to 44-year-olds is normally distributed with mean 125 in mm Hg and standard deviation 10 in mm Hg. Provide the interval for which 95% of the data will fall between.
Answer #1

μ = 125

σ = 10

Z value for 95% interval = Z0.025 = 1.96

upper limit of the interval = μ + 1.96 * σ = 125 + 1.96 * 10 = 144.6

lower limit of the interval = μ - 1.96 * σ = 125 - 1.96 * 10 = 105.4

The interval will fall between (105.4 , 144.6)